# Broadband localization of light at the termination of a topological photonic waveguide

### v2

@@
Abstract
@@
Manipulation of light in topological photonic platforms offers an exciting paradigm for exploring fundamental principles of light-matter interactions and developing novel communication technology components with new or improved performance [1, 2]. Specifically, photonic systems with non-trivial topology of their bulk eigenstates feature useful phenomena such as helical edge states [3, 4, 5, 6, 7, 8, 9, 10]. Careful tailoring of these photonic systems offers opportunities for efficient and robust molding of light on a nanophotonic chip, important for, e.g., coupling of quantum emitters [11], integrated photonic circuits [12, 13], and microcavity lasing [14, 15, 16].
@@
Topological photonic edge states can emerge at an interface of dielectric photonic crystals (PhCs) with broken spatial symmetry [4, 5]. Such broken symmetry can open a broad band gap in which bidirectional spin- or valley-locked chiral edge states exist. These edge states are known to exhibit suppressed backscattering by spin- or valley-flipping processes at corners that obey the protecting symmetry of the crystal [9, 17]. This notion of reduced backscattering and robust transmission raises the intriguing question of what happens to optical energy in a system if a topological waveguide is abruptly terminated. Related to this question, in nonreciprocal waveguides it was proposed that optical intensity could be enhanced at a waveguide termination, specifically in waveguides based on magneto-optic surface plasmon-polaritons [18, 19, 20, 21, 22, 23, 24] and on bi-anisotropic metamaterials [22]. These envisioned effects have been theorized to have topological origin [22] and it has been argued that their large and broadband field enhancements can benefit light-matter interactions and nonlinear optics [25, 26]. This suggests that reciprocal topological waveguides could also facilitate such enhancement, potentially broadening the scope of applications. Indeed, it was predicted that an abruptly terminated photonic valley Hall channel can localize light in a hotspot at the termination [27]. This topology-enabled buildup of optical energy is related to the near-conservation of the valley degree of freedom, as the minimization of intervalley overlap delays backreflection.
@@
In this work, we experimentally demonstrate localization and enhancement of optical energy at a termination of a topological waveguide at telecom frequencies, using near-field microscopy [28]. We observe that light intensity is enhanced by an order of magnitude for a broad range of frequencies where there exist no other surface modes that can carry energy away from the termination along the interface terminating the topological waveguide. We reveal that the symmetry of the termination plays an important role in the observed effect, confirming the responsible mechanism of backscattering reduction. These results thus demonstrate a new mechanism for electromagnetic field enhancement at the nanoscale, with consequently broad application potential in the control of light-matter interactions and improved manipulation and sensing of light in photonic technology.
@@
The topological waveguide is designed by connecting two mirror-inverted valley photonic crystals (VPCs) [5]. The VPCs have a rhombic unit cell that is composed of two opposite-facing triangular air holes with an inequivalence in size which breaks inversion symmetry and therefore opens a two-dimensional photonic band gap [5, 7, 29]. The inversion symmetry breaking results in an opposite-signed Berry curvature locally concentrated at the K/K’ valleys in the Brillouin zone. Connecting two VPCs that have distinct topologies, a valley-dependent edge state emerges which emulates the quantum valley Hall effect and is protected against intervalley scattering at defects that conserve the crystalline symmetry [8, 17]. Reciprocity guarantees that two edge states associated with opposite valley degree of freedom propagate in opposite directions. As long as the edge state dispersion lies below the light line, light is confined to the dielectric photonic crystal slab.
@@
The topological waveguide is abruptly terminated by an insulating trivial PhC with a two-dimensional photonic band gap that prohibits further propagation of the edge state. This termination results in a triple-junction of three inequivalent interfaces as shown in Fig. 1. The two interfaces between the VPCs and the PhC can in principle support trivial surface states. These states are not expected to span the entire VPC gap, as the difference in the valley Chern numbers of the VPC and the PhC is 1/2. Instead, each interface exhibits a frequency gap where no propagating states exist. We consider of particular interest the regime where the frequency gaps of both interfaces overlap. This frequency regime is bound by surface modes of both interfaces, and we hence refer to it as the surface mode gap. Operating at a frequency in this gap, light that is incident from the valley Hall waveguide towards the termination can only scatter out-of-plane (to free space radiation), or back into the topological waveguide. However, backscattering requires a flip of the valley index, which is suppressed at a suitably designed termination [27]. We investigate the resulting signature of suppressed backscattering on broadband optical enhancement at the termination.
@@ Results
When a valley edge mode encounters the termination, its evanescent tails penetrate into the two VPC bulks and are affected by the symmetry of the termination. As this can affect the rate of backscattering, we investigate the correlation between the observed localization and the symmetry of the termination. Along terminations with different geometries, the overlap between the forward propagating (K’-valley) and the backscattered (K-valley) modes are significantly different. Some unique choices of the termination can greatly suppress this overlap and hence reduce the rate of backscattering (see Supplementary Materials for detailed theoretical derivation). We distinguish between zigzag terminations, corresponding to or angles between the VPC-PhC and VPC-VPC interfaces, and armchair terminations, corresponding to angle. We reveal that zigzag terminations largely suppress the overlap, while armchair terminations feature significant overlap. This mathematical conclusion can be intuitively understood as following: a zigzag termination maximally follows the symmetry of the VPC lattice and thus conserves the valley DOF to a great degree, exploiting the topologically non-trivial nature of the VPC crystals. In contrast, an armchair termination strongly breaks the crystal symmetry that protects the VPC-VPC states from scattering efficiently. While a certain degree of symmetry breaking is unavoidable for any shape or termination, and light can thus eventually scatter back, the rate at which this happens thus depends strongly on the symmetry of the termination in the topological crystal [27]. Finally, we note that the finite size of the termination in the three-dimensional slab setting means that some scattering to free space is possible. There is in principle no reason to assume that the amount of free-space scattering as a result of this defect is significantly different for zigzag and armchair terminations.
@@
To study the effect of termination geometry experimentally, we probe the electromagnetic near-field intensity near the end of the waveguide for samples with either zigzag or armchair terminations. First, we experimentally confirm the optical energy localization at a zigzag termination. The green curve in Fig. 3A shows the intensity as a function of position along the VPC-VPC waveguide, which we obtain by integrating the intensity in a 2D scan over a transversal width of two lattice constants across the center of the waveguide. We observe for both the zigzag termination in Fig. 3A and the armchair termination in Fig. 3B an oscillatory behaviour in intensity along the majority of the VPC-VPC waveguide. This is the result of interference between forward and backward propagating modes, for all Bloch harmonics. Our phase-sensitive heterodyne detection scheme allows separating the intensities of forward and backward waves by applying a wavevector filter on the Fourier-transformed fields around the respective Bloch components [17]. The inverse Fourier transform using only the components that correspond to the forward (backward) propagating mode then returns the individual contribution to the intensity by that mode, shown by the red (blue) curve in a translationally invariant part of the topological waveguide. We refer to Supplementary Materials section 4 for a detailed explanation of the formalism behind the forward/backward separation. We observe a significant intensity of backward propagation for both the zigzag and armchair terminations. As the observation frequency resides in the surface mode gap, see Fig. 2D and Supplementary Materials Fig. S3D, there can be no energy propagation along the VPC-PhC interfaces. Light does, however, significantly scatter out-of-plane as a consequence of symmetry breaking at the termination, explaining a reduced intensity of the backward propagating mode. Furthermore, the intensity after being reflected from a zigzag or armchair termination is approximately the same, see Supplementary Materials Fig. S4B, indicating that the out-coupled energy to free space does not depend strongly on the type of termination.
@@
Importantly, the local intensity at the zigzag termination is observed to be 15 times larger than the average intensity of the forward propagating mode in the waveguide, in a sharply localized hot spot. In stark contrast, the armchair termination shows no enhancement, meaning that all energy is immediately reflected back. Simulations confirm that optical energy is strongly enhanced only at the zigzag termination, see Supplementary Materials Fig. S5. These observations are consistent with the theoretical finding that the local buildup of optical energy is strongly dependent on the termination orientation.
@@
The previous section studied the localized field enhancement for a particular laser frequency of 200.0 THz within the surface mode gap. We next investigate the enhancement in a broad frequency range. Fig. 4 shows a map of intensity as a function laser frequency and position along the waveguide. Each horizontal line is obtained in a similar way as the green integrated intensity line of Fig. 3. Three different terminations with varying dislocations are compared: two with a zigzag symmetry and one with an armchair symmetry. This comparison confirms that enhancement of field intensity only occurs at a zigzag termination and not at an armchair termination. We observe that the enhancement is present over 6 THz, and is the largest for frequencies within the surface mode gap. Its magnitude, which is the ratio between the maximum intensity of light at the termination and the average intensity of forward propagating light in the waveguide, varies with frequency, and reaches a maximum just below the upper limit of the gap (white dashed line in Fig. 4A). Supplementary Materials Fig. S4A depicts the spectral shape of this broadband peak.
@@
A possible interpretation of both the frequency dependence and the sharply defined spatial extent of the localized enhancement could be sought in the involvement of evanescent states of the VPC-PhC interface at the termination. Within the surface mode gap, the incident edge state cannot excite propagating states with a real wavevector along the VPC-PhC interfaces. However, a continuum of evanescent states with an imaginary wavevector along the interfaces can be expected to exist in the gap [30], which could be excited given appropriate wavefunction overlap with the incident waveguide field. These states cannot propagate, but could radiate into either free space or the backward propagating mode. In this interpretation, the suppressed scattering could be thought of as being assisted by compatible evanescent states at the interfaces, confined to the termination, that contribute to the buildup of local optical intensity there. As the overlap of incident and evanescent states would depend on the intricate details of their local fields, which are generally frequency dependent, their excitation amplitude can therefore vary slowly with frequency. In principle, evanescent modes would also exist at the armchair terminated waveguide with an overlap that may be significantly different from a zigzag terminated waveguide. The absence of any localization suggests that the intervalley scattering is much stronger and the edge states instead couple directly to the backward propagating mode. The significant difference in localization at this different symmetry hints that the coupling to evanescent states is negligible. This is indeed a remarkable distinction in scattering behaviour between the zigzag and armchair terminated waveguides. We note that this possible viewpoint of storage in local evanescent states does not provide an alternative explanation to the general effect of localization at a topological waveguide termination that we report, but it could provide a route to understanding the details of spatial and spectral extent of the intensity enhancement in future work.
@@ Discussion
This experimental observation of broadband light enhancement at a termination of a topological photonic waveguide is an example of a broader class of systems that could facilitate field enhancement due to suppressed backscattering [26]. Topologically non-trivial waveguides with a termination protected by lattice symmetry pave the way for the development of novel devices exploiting the potential of enhanced light matter interactions. Further research should explore the possible magnitude of broadband enhancement and the role of local design features, such as the dislocation and the choice of the terminating crystal. More broadly, the demonstrated paradigm of optical energy storage and localization, in combination with robust guiding and manipulation of light, present opportunities to on-chip photonic technology.
Acknowledgements

### v3

@@
Broadband localization of light at the termination of a topological photonic waveguide
@@ Introduction
Manipulation of light in topological photonic platforms offers an exciting paradigm for exploring fundamental principles of light-matter interactions and developing novel communication technology components with new or improved performance (?, ?). Specifically, photonic systems with non-trivial topology of their bulk eigenstates feature useful phenomena such as helical edge states (?, ?, ?, ?, ?, ?, ?, ?). Careful tailoring of these photonic systems offers opportunities for efficient and robust molding of light on a nanophotonic chip, important for, e.g., coupling of quantum emitters (?), integrated photonic circuits (?, ?), and microcavity lasing (?, ?, ?).
@@
Topological photonic edge states can emerge at an interface of dielectric photonic crystals (PhCs) with broken spatial symmetry (?, ?). Such broken symmetry can open a broad band gap in which bidirectional spin- or valley-locked chiral edge states exist. These edge states are known to exhibit suppressed backscattering by spin- or valley-flipping processes at corners that obey the protecting symmetry of the crystal (?, ?). This notion of reduced backscattering and robust transmission raises the intriguing question of what happens to optical energy in a system if a topological waveguide is abruptly terminated. Related to this question, in nonreciprocal waveguides it was proposed that optical intensity could be enhanced at a waveguide termination, specifically in waveguides based on magneto-optic surface plasmon-polaritons (?, ?, ?, ?, ?, ?, ?) and on bi-anisotropic metamaterials (?). These envisioned effects have been theorized to have topological origin (?) and it has been argued that their large and broadband field enhancements can benefit light-matter interactions and nonlinear optics (?). This suggests that reciprocal topological waveguides could also facilitate such enhancement, potentially broadening the scope of applications. Indeed, it was predicted that an abruptly terminated photonic valley Hall channel can localize light in a hotspot at the termination (?). This topology-enabled buildup of optical energy is related to the near-conservation of the valley degree of freedom, as the minimization of intervalley overlap delays backreflection.
@@
In this work, we experimentally demonstrate localization and enhancement of optical energy at a termination of a topological waveguide at telecom frequencies, using near-field microscopy (?). We observe that light intensity is enhanced by an order of magnitude for a broad range of frequencies where there exist no other surface modes that can carry energy away from the termination along the interface terminating the topological waveguide. We reveal that the symmetry of the termination plays an important role in the observed effect, confirming the responsible mechanism of backscattering reduction. These results thus demonstrate a new mechanism for electromagnetic field enhancement at the nanoscale, with consequently broad application potential in the control of light-matter interactions and improved manipulation and sensing of light in photonic technology.
@@
The topological waveguide is designed by connecting two mirror-inverted valley photonic crystals (VPCs) (?). The VPCs have a rhombic unit cell that is composed of two opposite-facing triangular air holes with an inequivalence in size which breaks inversion symmetry and therefore opens a two-dimensional photonic band gap (?, ?, ?). The inversion symmetry breaking results in an opposite-signed Berry curvature locally concentrated at the K/K’ valleys in the Brillouin zone. Connecting two VPCs that have distinct topologies, a valley-dependent edge state emerges which emulates the quantum valley Hall effect and is protected against intervalley scattering at defects that conserve the crystalline symmetry (?, ?). Reciprocity guarantees that two edge states associated with opposite valley degree of freedom propagate in opposite directions. As long as the edge state dispersion lies below the light line, light is confined to the dielectric photonic crystal slab.
@@
The topological waveguide is abruptly terminated by an insulating trivial PhC with a two-dimensional photonic band gap that prohibits further propagation of the edge state. This termination results in a triple-junction of three inequivalent interfaces as shown in Fig. 1. The two interfaces between the VPCs and the PhC can in principle support trivial surface states. These states are not expected to span the entire VPC gap, as the difference in the valley Chern numbers of the VPC and the PhC is 1/2. Instead, each interface exhibits a frequency gap where no propagating states exist. We consider of particular interest the regime where the frequency gaps of both interfaces overlap. This frequency regime is bound by surface modes of both interfaces, and we hence refer to it as the surface mode gap. Operating at a frequency in this gap, light that is incident from the valley Hall waveguide towards the termination can only scatter out-of-plane (to free space radiation), or back into the topological waveguide. However, backscattering requires a flip of the valley index, which is suppressed at a suitably designed termination (?). We investigate the resulting signature of suppressed backscattering on broadband optical enhancement at the termination.
@@ Results
When a valley edge mode encounters the termination, its evanescent tails penetrate into the two VPC bulks and are affected by the symmetry of the termination. As this can affect the rate of backscattering, we investigate the correlation between the observed localization and the symmetry of the termination. Along terminations with different geometries, the overlap between the forward propagating (K’-valley) and the backscattered (K-valley) modes are significantly different. Some unique choices of the termination can greatly suppress this overlap and hence reduce the rate of backscattering (see Supplementary Materials for detailed theoretical derivation). We distinguish between zigzag terminations, corresponding to or angles between the VPC-PhC and VPC-VPC interfaces, and armchair terminations, corresponding to angle. We reveal that zigzag terminations largely suppress the overlap, while armchair terminations feature significant overlap. This mathematical conclusion can be intuitively understood as following: a zigzag termination maximally follows the symmetry of the VPC lattice and thus conserves the valley DOF to a great degree, exploiting the topologically non-trivial nature of the VPC crystals. In contrast, an armchair termination strongly breaks the crystal symmetry that protects the VPC-VPC states from scattering efficiently. While a certain degree of symmetry breaking is unavoidable for any shape or termination, and light can thus eventually scatter back, the rate at which this happens thus depends strongly on the symmetry of the termination in the topological crystal (?). Finally, we note that the finite size of the termination in the three-dimensional slab setting means that some scattering to free space is possible. There is in principle no reason to assume that the amount of free-space scattering as a result of this defect is significantly different for zigzag and armchair terminations.
@@
To study the effect of termination geometry experimentally, we probe the electromagnetic near-field intensity near the end of the waveguide for samples with either zigzag or armchair terminations. First, we experimentally confirm the optical energy localization at a zigzag termination. The green curve in Fig. 3A shows the intensity as a function of position along the VPC-VPC waveguide, which we obtain by integrating the intensity in a 2D scan over a transversal width of two lattice constants across the center of the waveguide. We observe for both the zigzag termination in Fig. 3A and the armchair termination in Fig. 3B an oscillatory behaviour in intensity along the majority of the VPC-VPC waveguide. This is the result of interference between forward and backward propagating modes, for all Bloch harmonics. Our phase-sensitive heterodyne detection scheme allows separating the intensities of forward and backward waves by applying a wavevector filter on the Fourier-transformed fields around the respective Bloch components (?). The inverse Fourier transform using only the components that correspond to the forward (backward) propagating mode then returns the individual contribution to the intensity by that mode, shown by the red (blue) curve in a translationally invariant part of the topological waveguide. We refer to Supplementary Materials section 4 for a detailed explanation of the formalism behind the forward/backward separation. We observe a significant intensity of backward propagation for both the zigzag and armchair terminations. As the observation frequency resides in the surface mode gap, see Fig. 2D and Supplementary Materials Fig. S3, there can be no energy propagation along the VPC-PhC interfaces. Light does, however, significantly scatter out-of-plane as a consequence of symmetry breaking at the termination, explaining a reduced intensity of the backward propagating mode. Furthermore, the intensity after being reflected from a zigzag or armchair termination is approximately the same, see Supplementary Materials Fig. S4, indicating that the out-coupled energy to free space does not depend strongly on the type of termination.
@@
Importantly, the local intensity at the zigzag termination is observed to be 16 times larger than the average intensity of the forward propagating mode in the waveguide, in a sharply localized hot spot. In stark contrast, the armchair termination shows no enhancement, meaning that all energy is immediately reflected back. Simulations confirm that optical energy is strongly enhanced only at the zigzag termination, see Supplementary Materials Fig. S5. These observations are consistent with the theoretical finding that the local buildup of optical energy is strongly dependent on the termination orientation.
@@
The previous section studied the localized field enhancement for a particular laser frequency of 200.0 THz within the surface mode gap. We next investigate the enhancement in a broad frequency range. Fig. 4 shows a map of intensity as a function laser frequency and position along the waveguide. Each horizontal line is obtained in a similar way as the green integrated intensity line of Fig. 3. Three different terminations with varying dislocations are compared: two with a zigzag symmetry and one with an armchair symmetry. This comparison confirms that enhancement of field intensity only occurs at a zigzag termination and not at an armchair termination. We observe that the enhancement is present over 6 THz, and is the largest for frequencies within the surface mode gap. Its magnitude, which is the ratio between the maximum intensity of light at the termination and the average intensity of forward propagating light in the waveguide, varies with frequency, and reaches a maximum just below the upper limit of the gap (white dashed line in Fig. 4A). Supplementary Materials Fig. S4 depicts the spectral shape of this broadband peak.
@@
A possible interpretation of both the frequency dependence and the sharply defined spatial extent of the localized enhancement could be sought in the involvement of evanescent states of the VPC-PhC interface at the termination. Within the surface mode gap, the incident edge state cannot excite propagating states with a real wavevector along the VPC-PhC interfaces. However, a continuum of evanescent states with an imaginary wavevector along the interfaces can be expected to exist in the gap (?), which could be excited given appropriate wavefunction overlap with the incident waveguide field. These states cannot propagate, but could radiate into either free space or the backward propagating mode. In this interpretation, the suppressed scattering could be thought of as being assisted by compatible evanescent states at the interfaces, confined to the termination, that contribute to the buildup of local optical intensity there. As the overlap of incident and evanescent states would depend on the intricate details of their local fields, which are generally frequency dependent, their excitation amplitude can therefore vary slowly with frequency. In principle, evanescent modes would also exist at the armchair terminated waveguide with an overlap that may be significantly different from a zigzag terminated waveguide. The absence of any localization suggests that the intervalley scattering is much stronger and the edge states instead couple directly to the backward propagating mode. The significant difference in localization at this different symmetry hints that the coupling to evanescent states is negligible. This is indeed a remarkable distinction in scattering behaviour between the zigzag and armchair terminated waveguides. We note that this possible viewpoint of storage in local evanescent states does not provide an alternative explanation to the general effect of localization at a topological waveguide termination that we report, but it could provide a route to understanding the details of spatial and spectral extent of the intensity enhancement in future work.
@@ Discussion
This experimental observation of broadband light enhancement at a termination of a topological photonic waveguide is an example of a broader class of systems that could facilitate field enhancement due to suppressed backscattering (?). Topologically non-trivial waveguides with a termination protected by lattice symmetry pave the way for the development of novel devices exploiting the potential of enhanced light matter interactions. Further research should explore the possible magnitude of broadband enhancement and the role of local design features, such as the dislocation and the choice of the terminating crystal. More broadly, the demonstrated paradigm of optical energy storage and localization, in combination with robust guiding and manipulation of light, present opportunities to on-chip photonic technology.
Materials and Methods
Device fabrication
We fabricate a PhC slab on a SOI platform with a 220 nm thick silicon layer on a 3 µm buried oxide layer. First, a positive electron-beam resist of thickness 240 nm (AR-P 6200.09) is spin-coated. Then, the PhC design is patterned into the resist using electron-beam lithography on a Raith Voyager with 50 kV beam exposure. The electron-beam resist is developed in pentyl acetate/O-xylene/MIBK:IPA(9:1)/isopropanol, and the SOI chip subsequently undergoes reactive-ion etching in HBr and O2.
Next, a Suss MABA6 Mask Aligner patterns the photosensitive resist AZ1518 to define a selective wet-etching window on the PhC. After development with AZ400K:H2O, the buried-oxide layer is partially removed by wet etching in a 1:2 solution of 40% hydrofluoric acid and deionized water for 20 min. The sample is then subjected to critical point drying to obtain free-standing PhC membranes (?). Finally, to enable incoupling from the side, the SOI chip is cleaved on the end facets of the Si-ridge waveguides that extend to the PhCs. The Si-ridge waveguides are tapered down to support only the fundamental TE mode at the VPCs end facet. The ridge waveguides adiabatically transition to the VPC waveguide inside the lattice, enabling better index matching for efficient incoupling (?).
To account for fabrication imperfections that shift the operating frequencies of the devices, we fabricate triangular PhC lattices with three different lattice constants a0 = [449 nm, 494 nm, 539 nm]. In the VPCs, a rhombic unit cell contains two triangular holes of side lengths = 0.65a0 and = 0.38a0, respectively. A domain wall is created along VPC1 and VPC2 by applying a parity operation along the spatial y-coordinate. In the photonic bandgap material, a rhombic unit cell contains a single circular hole of radius r = 0.27a0.
Near-field microscopy setup
The near-field probe is fabricated by heated pulling of a single mode optical silica fiber using a P-2000 micropipette puller. Subsequently, the tips are etched in buffered oxide 7:1 for 20 minutes to ensure a wide opening angle of the aperture. The tips are coated homogeneously using a Temescal FC-20349 evaporator with first 1.5 nm chromium for improved adhesion and then 150 nm aluminium. The aluminium coating protects against radiative light loss at significant distances from the tip. The aluminium at the tip is then partly milled away by focused ion beam etching (Helios G4Cx) to create an aperture of 194 nm.
Probes are attached to one prong of a Farnell AB38T quartz tuning fork with a resonance of 32.768 kHz. The tuning fork has a Q-factor of 700 and is soldered to a custom made chip that drives the tuning fork, and reads out its oscillating frequency. The chip is then connected to a piezo dither block which rasters scan the probe above the photonic crystal while being controlled by a shear force feedback loop via a digital lock-in amplifier (Zurich Instruments HF2LI), keeping the tip at an expected height of 20 nm above the PhC surface.
Laser system
The light source deployed in the experiment is a continuous wave tunable laser (Santec TSL-710), which is operable in the infrared frequency range [1480 nm - 1640 nm]. In a heterodyne detection scheme, the beam is split into a signal and a reference path by a polarizing beam splitter and the frequency of the light in the reference path is shifted by 60 kHz using two acousto-optic modulators (AOMs). The signal path is then coupled into the photonic crystal ridge waveguide using a high numerical aperture (NA = 0.8) microscope objective. A near-field probe picks up a fraction of the evanescent fields above the sample. Subsequently, signal and reference path are recombined, which forms a beating signal at 60 kHz facilitating the extraction of the electric field components including phase resolution with a high signal-to-noise ratio via a pair of lock-in amplifiers (Stanford Research Systems SR830).
Finite Element Frequency-Domain Simulation
All numerical simulations of the photonic structure are performed using the finite element frequency-domain software COMSOL Multiphysics. Due to the high cost of 3D simulations on the computation resource, we simplify the simulation by leveraging the symmetry property of the structure. The suspended Si slab is mirror symmetric about the mid-height plane (, where t is the slab thickness). Therefore, all electromagnetic modes can be classified as TE-polarized [] and TM-polarized [] according to their field distributions at that plane (?). We focus on TE-polarized modes. By setting the mid-height plane of the slab as a perfect magnetic conductor, we simulate only half of the entire geometry, and only TE-polarized solutions are allowed. Above the slab (), an air domain () with a thickness of 3a0 is included in the simulation domain. Except for the mid-height plane of the slab, all other boundaries are set using the scattering boundary condition.
Acknowledgments
Acknowledgment